OpenStudy (anonymous):

When using the substitution method, what expression for x in the second equation would be substituted in the first equation? 4x - 7y = -9 x + 3y = 12

OpenStudy (anonymous):

hey think you can help me with a few others after this one

OpenStudy (campbell_st):

the easiest is x = 12 - 3y or you could use 3y = 12 - x ...... which can become y = 4 - x/3

OpenStudy (anonymous):

a. -3y + 12 b. 3y - 12 c. -3y - 12 d. 3y + 12

OpenStudy (anonymous):

these are the possible answers

OpenStudy (anonymous):

so b.

OpenStudy (anonymous):

is that right

OpenStudy (anonymous):

What is a possible first step in eliminating a variable in the following system of equations? x + 7y = -15 2x + 7y = -9 a. Multiply the first equation by -1. b. Multiply the second equation by -1 c. Multiply the first equation by -2. d. All of the above

OpenStudy (anonymous):

for the first question is the answer b